If a circle S ( x, y) = 0 touches at the point (2, 3) of the line x + y = 5 and (1,2) = 0, then radius of such circle.

Grade:12th pass

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find the answer to the your question below

Circle S passes through the point (2, 3).
(1, 2) is the centre of S.
Radius of circle:
r = \sqrt{(2-1)^{2}+(3-2)^{2}}
r = \sqrt{1+1}
r = \sqrt{2}